Prove or disprove: The average of three real numbers is greater than or equal to at least one of the numbers.


1
Expert's answer
2021-11-10T17:30:20-0500

Let us prove that the average of three real numbers is greater than or equal to at least one of the numbers.

Let "a_1,a_2,a_3" be arbitrary real numbers, and "b=\frac{a_1+a_2+a_3}3" be their average value. Let us prove by contraposition. Suppose that "b<a_1" and "b<a_2" and "b<a_3." Then "3b=b+b+b<a_1+a_2+a_3," and hence "b<\frac{a_+a_2+a_3}3=b." We get the contradiction "b<b." This contradictions prove that the average of three real numbers is greater than or equal to at least one of the numbers.